OpenStudy (twistrose):

2x-15/2=x-7 Identify whether the equation has zero, one, or infinitely many solutions. zero solutions one solution infinitely many solutions

OpenStudy (jango_in_dtown):

only one solution.. solve for x

OpenStudy (jango_in_dtown):

2x-x=15/2-7 i.e. x=0.5

OpenStudy (twistrose):

ok i got it thank you so much

OpenStudy (twistrose):

wait it said it is wrong

OpenStudy (twistrose):

nvm it was zero solutions thanks anyway

OpenStudy (jango_in_dtown):

Did you type the correct question? If yes, then my answer is correct.

OpenStudy (twistrose):

does this help

OpenStudy (jango_in_dtown):

see you didnot type correctly. 2x-15/2 is not the same as (2x-15)/2

OpenStudy (jango_in_dtown):

in the given problem it will be then x-7.5=x-7 i.e. 7.5=7, which is not true. thus zero solutions

OpenStudy (twistrose):

ok im sorry thank you i thought it was the same